OpenStudy (anonymous):

(will give medals) Question 4. 4. Solve the following system of equations by linear combination: 3d – e = 7 d + e = 5 (Points : 1) There is no solution. There are an infinite number of solutions. The solution is (3, 2). The solution is (2, –1).

OpenStudy (anonymous):

linear combination is the same as elimination. 3d - e = 7 d + e = 5 --->(-3) ---------------- 3d - e = 7 -3d - 3e = -15 (result of multiplying by -3) ----------------add -4e = - 8 -- divide both sides by -4 e = 2 now sub 2 in for e in either equation to find d 3d - e = 7 3d - 2 = 7 3d = 7 + 2 3d = 9 d = 3 solution : d = 3 and e = 2 or (3,2) any questions ?